# Autoassociative Learning of Structural Representations for Modeling and Classification in Medical Imaging

### v2

@@ Abstract
Deep learning architectures based on convolutional neural networks tend to rely on continuous, smooth features. While this characteristics provides significant robustness and proves useful in many real-world tasks, it is strikingly incompatible with the physical characteristic of the world, which, at the scale in which humans operate, comprises crisp objects, typically representing well-defined categories. This study proposes a class of neurosymbolic systems that learn by reconstructing the observed images in terms of visual primitives and are thus forced to form high-level, structural explanations of them. When applied to the task of diagnosing abnormalities in histological imaging, the method proved superior to a conventional deep learning architecture in terms of classification accuracy, while being more transparent.
@@ 1 Introduction
Annotation of medical imaging is notoriously time-consuming, prone to human biases, and hard to reconcile with the insatiable demands of contemporary machine learning. Deep Learning (DL) models trained on annotated data are often narrow in focusing on features that are specific to a given context (anomaly, pathology, etc.) rather than discovering and capturing general characteristics of observed structures and processes, which may make them susceptible to deceptive image features and lead to inferior generalization.
We posit that one of the primary causes of this challenge is the unstructured character of DL architectures. Contemporary DL models are essentially intertwined compositions of dot products and nonlinearities, conglomerates of often billions of unsophisticated units that process data in a highly distributed and continuous, non-symbolic fashion. Their training requires large volumes of data, which are often hard to come by, and involves exorbitant amounts of compute and energy. If the task is posed within the supervised learning paradigm, those data need to be not only curated, but also annotated (labeled), which limits their availability even further. Last but not least, as each processing unit takes care only of a minuscule fraction of inference, it is very hard to explain the model and its decisions to a human in a transparent and succinct fashion.
In this study, we argue for stronger involvement of unlabeled data in the construction of analytic and diagnostic ML models and propose ASR, a neurosymbolic architecture trained to form Auto-associative Structural Representations, in which a generative decoder synthesizes physically plausible structural models that explain the observed image. ASR combines a convolutional encoder queried on a raster image with a symbolic decoder that performs differentiable rendering to reproduce that image. As the involved symbolic components are differentiable, the entire architecture can be efficiently trained with conventional gradient-based algorithms. ASR can learn effectively from relatively small data sets, offers better interpretability than conventional models, and promises better diagnostic accuracy.
2 The proposed method
2.1 Motivations
The blueprint of convolutional neural networks (convnets), the working horse of DL applications in computer vision (CV), assumes that aggregation of input pixels gives rise to higher-level features formed at consecutive layers, which are expected to convey information about (among others) structure and can be then used for inference. While this scheme proved very useful and effective in multiple contexts, its structural aspect is only implicit in the sense of depending on the interpretation of weights in units’ receptive fields. In particular, convnets have no means of expressing the ‘objectness’ of their percepts. The entirety of processing is about manipulating pixels, whether these are literally the pixels available in the input image, or ‘pixels’ that embody higher-level features formed in further layers, usually at a lower spatial resolution.
The main motivation behind the proposed approach is to learn image interpretation in terms of representations that are explicitly more structural and hierarchical than in the mainstream DL, and in particular those learned by convnets. To this aim, we equip our model with the capacity of reasoning in terms of (relatively simple) visual primitives, rather than individual pixels. This representation is explicit and enforced, i.e. the model has to reason about the image in the prescribed terms.
To provide for the above enforcement, we resort to autoassociative learning: ASR is a neurosymbolic autoencoder that, given an input image, attempts to reproduce it with parameterized visual primitives. The reproduced image is compared to the input image using a reconstruction loss that is minimized with gradient in training.
We assume the visual primitives to be spatially continuous (so that they correspond to connected regions when rendered on rasters), compact, and tunable with parameters that allow the model to control their visual properties, like size, orientation, shape, and color (appearance). In this preliminary study, we resort to ellipses as a simple yet nontrivial class of primitives. We exemplify the approach by applying it to microscopic images of stained histological sections (Sec. 4); nevertheless, it can be generalized to other domains, including 3D volumetric imaging.
@@ 2.1 Motivations
2.2 The architecture and training
The ASR autoencoder comprises three main components: an Encoder, a number of Modelers working in parallel, and a Renderer (corresponding to the decoder in conventional autoencoder architectures) (Fig. 1).
@@ 2.2 The architecture and training
There is a separate modeler for each spatial scale $j$ (Fig. 1), realized as a trainable $1\times 1$ convolution layer that maps $z_{j}$ at each location in the image to six variables that control the rendering of the primitive at that location: the horizontal and vertical scaling factors ($w_{j}$ and $h_{j}$), the rotation angle $d_{j}$, and the three RGB components $a_{j}$ that control the primitive’s color appearance. The layer features the sigmoid activation function to bound the output interval. Subsequently, $w_{j}$ and $h_{j}$ are linearly scaled to $[0.1,2]$, so that it becomes impossible to ‘collapse’ the primitive to a point, while the upper bound of 2 allows partial overlap of primitives. $d_{j}$ is scaled to $[0,2\pi]$.
As argued later, the quality of reproduction that guides the training process of ASR is only a means to the end of learning informative structural representations. Therefore, the absolute values of that metric are not very relevant. For this reason, and to avoid excessive computational cost, we resort to spatially sparse rendering of visual primitives, rather than allowing them to be ‘anchored’ at any pixel in the image. We achieve this with two means. Firstly, as in many convnets, the dimensions of the spatial latent decrease with scale, so that the receptive fields of individual elements of those latents, expressed in terms of the input image, become larger with consecutive scales. Secondly, each Modeler is applied to its spatial latent with a stride greater than 1. As a result, the primitives at scale $j$ are produced at certain spacing $r_{j}$ and overlap only when the scaling factors $w_{j}$ and $h_{j}$ are relatively high. Technically, we set $r_{j}$ to half the size of the receptive field at scale $j$.
In particular, we will show in Sec. 4 that such a representation, though not very precise at reproduction, is very informative for solving a concrete classification task.
@@
As ASR is meant to be trained end-to-end with gradient, the rendering process must be differentiable. While prior work on differentiable rendering exists (see, e.g., (Ravi et al., 2020; Li et al., 2020; Fuji Tsang et al., 2022; Jakob et al., 2022)), the elliptical shapes of our primitives allow achieving this goal with a relatively crude rendering of a blurry ‘blobs’ rather than ‘crisp’ ellipses. We realize this for a single primitive ellipse at a given location at scale $j$ with the following steps:
@@
    Create a monochrome (single-channel) raster $R_{j}$ of size $(2r_{j}+1)\times(2r_{j}+1)$.
@@
    Apply $A$ to $R_{j}$: generate a transformed grid from $A$ and use it to sample from $R$ with bilinear interpolation (so this step is differentiable too).
@@
The $R_{j}$s for all grid cells are subsequently fused into a single canvas that represents the rendering outcome at the $j$th scale. To this aim, the individual $R$s are spatially co-registered, i.e. placed at grid cells in the coordinate system of the canvas, and their pixel values are aggregated. As in this study, we are interested in imaging in transmissive mode, where the light is absorbed by objects (rather than reflected or emitted by them), we perform multiplicative aggregation of the complements of $R$s in each RGB channel independently, so that the resultant channel value of a pixel $(x,y)$ in the canvas is $\Pi_{i}(1-R_{j}^{(i)}(x,y))$ where $R_{j}^{(i)}(x,y)$ is the channel value of pixel $(x,y)$ in $i$th raster at scale $j$, after spatial co-registration. [^1] As any given raster overlaps only with the neighboring rasters, this aggregation is limited to the neighborhood.
@@
It is worth noticing that the Renderer does not involve any trainable parameters, and therefore its only role in training is to back-propagate the feedback received from the reconstruction error (the loss function) to the preceding components of the architecture, viz. the Modeler and the Encoder.
Training. ASR is trained with gradient descent by minimizing the loss function defined as the mean square error between the reproduction produced by the Renderer and the input image, averaged over pixels and RGB channels. We ignore the image margin of width 16 pixels in this calculation, to lessen the impact of border effects (the model may be struggling to reproduce the structures located close to the borders, as they can be only partially visible, extending beyond the image). We refer to this cropped variant of MSE as Masked MSE (MMSE).
Preliminary experiments indicated the importance of parameter initialization, which is due to the unusual/nontrivial processing taking place in the Renderer and the presence of sigmoid functions in Modelers. Therefore, we initialize the weights of ConvBlocks and Modelers with the Xavier method (Glorot and Bengio, 2010), but sample their biases from $N(0,1)$ and set the bias of the BackgroundBlock to 1 to encourage it to participate in the rendering process.
@@ 3 Related work
Krawiec and Nowinowski (2024) recently proposed a related neurosymbolic architecture that uses a domain-specific language to capture selected priors of image formation, including object shape (represented using elliptic Fourier transform), appearance, categorization, and geometric transforms. They express template programs in that language and learn their parameterization with features extracted from the scene by a convnet. When executed, the parameterized program produces geometric primitives which are rendered and assessed for correspondence with the scene content and trained via auto-association with gradient. The authors confront their approach with a baseline method on a synthetic benchmark and demonstrate its capacity to disentangle selected aspects of the image formation process, learn from small data, correct inference in the presence of noise, and out-of-sample generalization.
@@ 4 Experimental validation
The experimental validation presented in the following comprises Stage 1 of representation learning (Sec. 4.4) and Stage 2 of learning to classify (Sec. 4.5). In the former, we train ASR and baseline models (all autoencoders) on a training set of images via autoassociation, to allow them to form informative latent representations. The primary objective in this stage is the quality of reconstruction. In Stage 2, we compose the encoders extracted from the models obtained in Stage 1 with decision trees to perform image classification. The main research question is whether the structure-aware representations formed by ASR in Stage 1 translate in Stage 2 to better test-set accuracy than the baseline models.
@@ 4.1 Problem statement and data
Figure 2: Example Thyroid Gland Whole Slide Image from the Biospecimen Research Database (BRD) Bio (2023)
As the challenges outlined in the Introduction are particularly relevant for medical imaging, we focus on this realm. Given that ASR, as described in Sec. 2, is designed to reconstruct images specifically with elliptical shapes, we choose microscopic histological imaging, where cells and other structures can be expected to lend themselves to modeling with such primitives. More specifically, we consider images of the human thyroid as the subject of this study. Thyroid follicles filled with colloid hormones form the major mass of the organ, and are usually oval, or almost round; see examples in Figs.2 and 3. Yet they vary in shapes, sizes, orientations, colors, and spatial arrangements, which makes them a perfect yardstick for the kind of modeling implemented by ASR.
@@
We rely on the Biospecimen Research Database (Bio, 2023) as the source of examinations for this study. This database is a large repository of digital slides containing, among others, 893 Whole Slide Images (WSIs) of the thyroid gland, each at least 50k $\times$ 50k pixels in size and typically covering a specimen of at least 10 by 10 mm (see example slide in Fig. 2). The slides are available in Aperio SVS format, which we parsed using the slideio library (Janowczyk, 2020) that allows reading in both image data and metadata.
From this collection, we selected a representative sample of 30 examinations representing three clinical conditions (decision classes): Benign (no abnormalities), Hashimoto, and Nodularity. When sampling, we provided balanced distributions of sex and age of patients (Table 3). The class assignment is based on the golden standard, as present in the clinical review comments included in the metadata. The selected sample is publicly available. [^2] https://drive.google.com/file/d/1xpjTIJ1pB9KawD7AveaVTq77OzpM8WwO/view?usp=share_link Example fragments of the slides are shown in Figure 3. Though dimensions of images vary significantly in the sample, their spatial resolution is the same (0.4942 $\mu$m/pix). The apparent magnification (AppMag) is also constant across the dataset (20).
We divide the 30 selected examinations into training set (15 examinations), validation set (6 examinations) and test set (9 examinations). This sampling was stratified, so that these proportions hold also in each decision class. Each examination comes from a different patient.
WSIs are too large to process them directly with DL models and, as exemplified in Fig. 2, usually contain a substantial fraction of white, uninformative background. Therefore, we extract patches from WSIs by scanning them systematically with a 1024$\times$1024 sliding window at horizontal and vertical stride of 635 pixels (2/3 of window size) and accept only the patches that are occupied with tissue in at least 80%. Then, we downsize them to $256\times 256$ pixels. This results in from 400 to 1550 256$\times$256 patches per WSI. As a result, the train:validation:test proportions of 15:6:9 on the level of examinations translate to 10,915:4,945:7,235 on the level of image patches. As a side-effect, the distributions of decision classes in these subsets are not uniform and diverge significantly (Fig. 4).
@@ 4.2 Configuration of ASR
We configure ASR so that its encoder comprises 3 ConvBlocks (Sec. 2.2), which implies that modeling and reconstruction take place at 3 scales/resolutions corresponding to those blocks: the highest (finest, $j=2$, grid spacing $r_{2}=128$), the intermediate ($j=1$, $r_{1}=64$), and the lowest (coarsest, $j=0$, $r_{0}=32$). The grid dimensions at those scales are, respectively, 2$\times$2 cells, 4$\times$4 cells, and 8$\times$8 cells. This implies that the encoder parameterizes 84 ellipses in total. As each of them requires 6 parameters, the total dimensionality of the structural latent is 504.
All models are designed to process the downsized $256\times 256$ patches. They were trained for up to the number of epochs indicated in the following, unless the PyTorch’s EarlyStopping component signaled stagnation on the validation set (with patience of 20 epochs). For details on other common settings and elements of the configuration, see Section B.
@@ 4.3 Compared variants of the method
Regularized. In preliminary experimenting, ASR exhibited a significant tendency to rely primarily on rendering at higher resolutions (on the denser grids), while neglecting the lowest resolutions. As this is at odds with our goal of learning succinct and transparent representations, we devise a regularization mechanism to curb the total number of graphical primitives used in rendering. We define the Appearance Regularization Value as:
[…]
where $\widehat{y}$ is the reconstruction image, $J$ is the set of all scales, $Loc_{J}$ is the set of variables describing shapes for all scales in $\widehat{y}$, $Loc_{j}$ is the set of variables describing shapes at scale $j$, $w_{j}$ is the weight of scale $j$, $a^{j}_{l}$ is the three-dimensional vector of RGB color components at scale $j$ and location $l$, and $\alpha$ is the factorization parameter $\in(0.5,1.0)$. Based on preliminary experimenting, we set the scale weights as follows: $w_{0}=0.6$, $w_{1}=0.9$, and $w_{2}=1.2$. The trade-off between this regularization term and the quality of reconstruction is controlled by weighing ARV in the total loss:
$Loss(y,\widehat{y})=\textrm{MMSE}(y,\widehat{y})+\lambda_{a}ARV(\widehat{y}),$ (2)
@@
Incremental. While the models trained with regularization indeed turned out to use high-resolution rendering more sparingly than the base configurations, we sought other means of addressing their over-representation. We devised an incremental training procedure, in which the model is forced to rely more on the coarser resolutions earlier in training, while with time being gradually allowed to use the higher-resolution primitives. This is achieved by multiplying all outputs of the Modeler for scale $j$ by $\beta_{j}$. We start with $\beta_{0}=1$ and $\beta_{1}=\beta_{2}=0.01$. Then, $\beta_{j}$s are increased after each epoch according to the formula:
$\mathit{\beta_{j}\leftarrow\min(1.0,\beta_{j}+\gamma_{j}).}$ (3)
Based on preliminary experimenting, we set $\gamma_{1}=\gamma_{2}=0.1$ and $\gamma_{0}=0$, i.e. the rendering at the lowest-resolution scale is unaffected by this mechanism and remains constant throughout training. Importantly, the above incrementing commences only in epoch 8 for scale 1 and in epoch 20 for scale 2.
@@
Baseline. As a baseline, we use a conventional convolutional autoencoder with a global latent. Its encoder starts with a stack of encoder blocks, each comprising a convolution layer, an Exponential Linear Unit (ELU) activation function, batch normalization, and a 2D max pooling layer. This is followed by flattening to a vector and a single dense layer that produces the global latent. We set the dimensionality of the global latent to 200 as a compromise between the computational cost and reconstruction quality. The decoder is a stack of decoder blocks, each comprising a 2D up-sampling layer with nearest neighbor interpolation, a convolution layer with ELU activation function, and batch normalization. As for ASR, the input and output are 256$\times$256 RGB images. The architecture has $3{,}781{,}509$ trainable parameters and has been trained for up to 50 epochs.

### v3

@@ Abstract
Deep learning architectures based on convolutional neural networks tend to rely on continuous, smooth features. While this characteristics provides significant robustness and proves useful in many real-world tasks, it is strikingly incompatible with the physical characteristic of the world, which, at the scale in which humans operate, comprises crisp objects, typically representing well-defined categories. This study proposes a class of neurosymbolic systems that learn by reconstructing images in terms of visual primitives and are thus forced to form high-level, structural explanations of them. When applied to the task of diagnosing abnormalities in histological imaging, the method proved superior to a conventional deep learning architecture in terms of classification accuracy, while being more transparent.
@@ 1 Introduction
The blueprint of convolutional neural networks (convnets), the working horse of deep learning (DL) applications in computer vision (CV), assumes that raster-based processing at consecutive layers gives rise to higher-level features that convey information about objects, the structures they form, and other aspects of scene content. While this proved effective in many use cases, the structural reasoning in convnets is only implicit in the sense of residing in weights in units’ receptive fields. In particular, convnets have no means of explicitly capturing the ‘objectness’ of percepts. This is strikingly at odds with the physical characteristic of the world, which, at the scale in which humans operate, is full of objects that have well-defined properties such as shape, size, orientation, and color.
This incompatibility with the characteristics of natural images has several downsides. Firstly, it increases the risk of overfitting, as the combinatorial capacity of raster-based processing is excessively expressive for capturing natural scenes. To curb this risk, convnets need to be trained on huge volumes of data, which can be particularly troublesome in supervised settings that require image annotation. Last but not least, they offer little explanatory capability and need specialized algorithms to elucidate the processing or justify the decisions being made.
To address the above challenges, we propose ASR, a neurosymbolic autoencoder that forms Auto-associative Structural Representations, physically plausible scene descriptions that capture and explain the observed image in terms of visual primitives rather than individual pixels. This representation is explicit, i.e. the model is forced to reason in the prescribed terms by autoassociative learning: given an input image, ASR attempts to reproduce it by parameterizing the primitives. The architecture comprises a convolutional encoder queried on a raster image and a symbolic decoder that performs differentiable rendering of primitives to reproduce that image. As all these components are differentiable, the architecture can be efficiently trained end-to-end with conventional gradient-based algorithms, using a simple pixel-wise reconstruction loss that compares the reproduced image with the input image.
ASR can work with any visual primitives that can be rendered in a differentiable fashion. However, to provide alignment with the characteristics of natural images and to improve explainability, it is desirable for the primitives to be spatially continuous, compact, and tunable with parameters that independently control human-interpretable visual properties, such as size, orientation, shape, and color. In this study, we resort to ellipses as a simple yet nontrivial class of primitives. We demonstrate the approach by applying it to microscopic images of stained histological sections (Sec. 4). The ASR is first trained on the (unannotated) images via autoassociation. Then, we use the learned representation to build a transparent diagnostic model (a decision tree), demonstrate its predictive superiority to a baseline configuration, and show how the decisions it makes can be explained with the interpretable properties of visual primitives.
@@
2 ASR architecture and training
The ASR autoencoder comprises an Encoder, a number of Modelers, and a Renderer – the last one corresponding to the decoder in conventional autoencoders (Fig. 1).
@@ 2 ASR architecture and training
There is a separate modeler for each spatial scale $j$ (Fig. 1), realized as a trainable $1\times 1$ convolution layer that maps $z_{j}$ at each location in the image to six variables that control the rendering of the primitive at that location: the horizontal and vertical scaling factors ($w_{j}$ and $h_{j}$), the rotation angle $d_{j}$, and the tuple of RGB components $a_{j}$ that control the primitive’s color appearance. The layer features the sigmoid activation function to bound the output intervals. Subsequently, $w_{j}$ and $h_{j}$ are linearly scaled to $[0.1,2]$, so that it becomes impossible to ‘collapse’ the primitive to a point, while the upper bound of 2 allows partial overlap of primitives. $d_{j}$ is scaled to $[0,2\pi]$.
As argued later, the quality of reproduction that guides the training process of ASR is only a means to an end of learning informative structural representations. Therefore, the absolute values of that metric are not very relevant. For this reason, and to avoid excessive computational cost, we resort to spatially sparse rendering of visual primitives, rather than allowing them to be ‘anchored’ at any pixel in the image. We achieve this with two means. Firstly, as in many convnets, the dimensions of the spatial latent decrease with scale, so that the receptive fields of individual elements of those latents, expressed in terms of the input image, become larger with consecutive scales. Secondly, each Modeler is applied to its spatial latent with a stride greater than 1. As a result, the primitives at scale $j$ are produced at certain spacing $r_{j}$ and overlap only when the scaling factors $w_{j}$ and $h_{j}$ are relatively high. Technically, we set $r_{j}$ to half the size of the receptive field at scale $j$.
@@
As ASR is meant to be trained end-to-end with gradient, the rendering process must be differentiable. While prior work on differentiable rendering exists (see, e.g., (Ravi et al., 2020; Li et al., 2020; Fuji Tsang et al., 2022; Jakob et al., 2022)), the elliptical shapes of our primitives allow achieving this goal with a relatively crude rendering of a blurry ‘blobs’ rather than ‘crisp’ ellipses. For a given ellipse at scale $j$, we realize this with the following steps:
@@
    Create a monochrome (single-channel) zeroed raster $R_{j}$ of size $(2r_{j}+1)\times(2r_{j}+1)$.
@@
    Apply $A$ to $R_{j}$: generate a transformed grid of points from $A$ and use it to sample from $R$ with bilinear interpolation (so this step is differentiable too).
@@
The $R_{j}$s for all grid cells are subsequently fused into a single canvas that represents the rendering outcome at the $j$th scale. To this aim, the individual $R$s are spatially co-registered, i.e. placed at grid cells in the coordinate system of the canvas, and their pixel values are aggregated. As in this study, we are interested in imaging in transmissive mode, where the light is absorbed by objects (rather than reflected or emitted by them), we perform multiplicative aggregation of the complements of $R$s in each RGB channel independently, so that the resultant channel value of a pixel $(x,y)$ in the canvas is $\Pi_{i}(1-R_{j}^{(i)}(x,y))$ where $R_{j}^{(i)}(x,y)$ is the channel value of pixel $(x,y)$ in $i$th raster at scale $j$, after spatial co-registration.
@@
Notice that the Renderer does not involve any trainable parameters.
Training. ASR is trained with gradient descent by minimizing the loss function defined as the mean square error between the reproduction produced by the Renderer and the input image, averaged over pixels and RGB channels. We ignore the margin of 16 pixels in this calculation, which translates to approximately 6% of the image width, to lessen the impact of border effects (the model may be struggling to reproduce the partially visible structures that extend beyond the image). We refer to this cropped variant of MSE as Masked MSE (MMSE).
Preliminary experiments indicated the importance of parameter initialization, which is due to the highly nonlinear processing taking place in the Renderer and the presence of sigmoid functions in Modelers. Therefore, we initialize the weights of ConvBlocks and Modelers with the Xavier method (Glorot and Bengio, 2010), but sample their biases from $N(0,1)$ and set the bias of the BackgroundBlock to 1 to encourage it to participate in the rendering process.
@@ 3 Related work
Krawiec and Nowinowski (2024) recently proposed a related neurosymbolic architecture that uses a domain-specific language to capture selected priors of image formation, including object shape (represented by elliptic Fourier transform), appearance, categorization, and geometric transforms. They express template programs in that language and learn their parameterization with features extracted from the scene by a convnet. When executed, the parameterized program produces geometric primitives which are rendered in a differentiable fashion.
@@ 4 Experimental validation
The experimental validation presented in the following comprises representation learning (Stage 1, Sec. 4.4) and a ‘downstream’ task of learning to classify (Stage 2, Sec. 4.5). In the former, we train ASR and baseline models (all autoencoders) on a training set of images via autoassociation, to allow them to form informative latent representations. In Stage 2, we compose the Encoders and Modelers extracted from the models obtained in Stage 1 with decision trees to perform image classification. The main research question is whether the structure-aware representations formed by ASR in Stage 1 translate in Stage 2 to better test-set accuracy than the baseline models.
@@ 4.1 Problem statement and data
As the challenges outlined in the Introduction (data inefficiency, costly annotation, limited explainability) are particularly relevant for medical imaging, we focus on this realm. Given that ASR, as described in Sec. 2, is designed to reconstruct images specifically with elliptical shapes, we choose microscopic histological imaging, where cells and other structures can be expected to lend themselves to modeling with such primitives. More specifically, we consider images of the human thyroid as the subject of this study. Thyroid follicles filled with colloid hormones form the major mass of the organ, and are usually oval, or almost round; see examples in Figs.  2 and 5. Yet they vary in shapes, sizes, orientations, colors, and spatial arrangements, which makes them a perfect yardstick for the kind of modeling implemented by ASR.
@@
We rely on the Biospecimen Research Database (BRD, ) as the source of examinations for this study (Sec. A). From this collection, we selected a sample of 30 examinations representing three clinical conditions (decision classes): Benign (no abnormalities), Hashimoto, and Nodularity (Fig. 2). When sampling, we provided balanced distributions of sex and age of patients (Table 3). The class assignment is based on the golden standard, as present in the clinical review comments included in the metadata.
We divide the 30 selected examinations into the training set (15 examinations), the validation set (6 examinations) and the test set (9 examinations). This partitioning was stratified, so that these proportions hold also in each decision class. Each examination comes from a different patient.
WSIs are too large to process them directly with DL models and, as exemplified in Fig. 5, usually contain a substantial fraction of white, uninformative background. Therefore, we extract patches from WSIs by scanning them systematically with a 1024$\times$1024 sliding window at horizontal and vertical stride of 635 pixels (2/3 of window size) and accept only the patches that are occupied with tissue in at least 80%. Then, we downsize them to $256\times 256$ pixels. This results in from 400 to 1550 256$\times$256 patches per WSI. As a result, the train:validation:test proportions of 15:6:9 on the level of examinations translate to $10{,}915$:$4{,}945$:$7{,}235$ (i.e. approximately 15:6.8:9.9) on the level of image patches. As a side-effect, the distributions of decision classes in these subsets are not uniform and diverge significantly (Fig. 6).
@@ 4.2 Configuration of ASR
We configure ASR so that its encoder yields latent from three ConvBlocks (Sec. 2) and modeling and reconstruction take place at scales/resolutions corresponding to those blocks: the highest (finest, $j=2$, grid spacing $r_{2}=128$), the intermediate ($j=1$, $r_{1}=64$), and the lowest (coarsest, $j=0$, $r_{0}=32$). The grid dimensions at those scales are, respectively, 2$\times$2 cells, 4$\times$4 cells, and 8$\times$8 cells, so that the encoder parameterizes 84 ellipses in total. As each of them requires 6 parameters, the overall dimensionality of the structural latent amounts to 504. All models are designed to process the downsized $256\times 256$ patches and trained until the PyTorch’s EarlyStopping component signaled stagnation on the validation set (with patience of 20 epochs), for the number of epochs listed in B alongside other settings.
@@ 4.3 Compared variants of the method
Regularized. In preliminary experimenting, ASR exhibited a significant tendency to rely primarily on rendering at higher resolutions (on the denser grids), while neglecting the lowest resolutions. As this is at odds with our goal of learning succinct and transparent representations, we devise the Appearance Regularization Value to curb the number of graphical primitives used in rendering:
$ARV(\hat{y})=\frac{1}{|Loc_{J}|}\sum_{j\in J}w_{j}\sum_{l\in Loc_{j}}\sum_{c\in{a_{j}^{l}}}c^{\alpha},$ (1)
where $\hat{y}$ is the reconstruction image, $J$ is the set of all scales, $Loc_{J}$ is the set of variables describing shapes for all scales in $\hat{y}$, $Loc_{j}$ is the set of variables describing shapes at scale $j$, $w_{j}$ is the weight of scale $j$, $c$ is the value corresponding to RGB color descriptors at scale $j$ and location $l$, and $\alpha$ is the factorization parameter $\in(0.5,1.0)$. Based on preliminary experimenting, we set the scale weights as follows: $w_{0}=0.6$, $w_{1}=0.9$, and $w_{2}=1.2$. The trade-off between this regularization term and the quality of reconstruction is controlled by weighing ARV in the total loss:
$Loss(y,\hat{y})=\textrm{MMSE}(y,\hat{y})+\lambda_{a}ARV(\hat{y}),$ (2)
@@
Incremental. While the Regularized models indeed turned out to render at high resolution more sparingly than the Base configurations, we sought other means of addressing their over-representation using incremental training, in which the model is forced to rely more on the coarser resolutions earlier in training, while with time being gradually allowed to use the higher-resolution primitives. This is achieved by multiplying all outputs of the Modeler for scale $j$ by $\beta_{j}$. We start with $\beta_{0}=1$ and $\beta_{1}=\beta_{2}=0.01$. Then, $\beta_{j}$s are increased after each epoch according to the formula $\mathit{\beta_{j}\leftarrow\min(1.0,\beta_{j}+\gamma_{j}).}$ Based on preliminary experimenting, we set $\gamma_{1}=\gamma_{2}=0.1$ and $\gamma_{0}=0$, i.e. the rendering at the lowest resolution is unaffected by this mechanism and remains constant throughout training. The incrementing commences in epoch 8 for scale 1 and in epoch 20 for scale 2.
@@
Baseline. As a baseline, we use a conventional convolutional autoencoder with a global latent. Its encoder starts with a stack of encoder blocks, each comprising a convolutional layer, an Exponential Linear Unit (ELU) activation function, batch normalization, and a 2D max pooling layer. This is followed by flattening to a vector and a single dense layer that produces the global latent. We set the dimensionality of the global latent to 200 as a compromise between the computational cost and reconstruction quality. The decoder is a stack of decoder blocks, each comprising a 2D up-sampling layer with nearest neighbor interpolation, a convolutional layer with ELU activation function, and batch normalization. As for ASR, the input and output are 256$\times$256 RGB images. The architecture has $3{,}781{,}509$ trainable parameters and has been trained for up to 50 epochs.